OpenStudy (anonymous):

Help? 1. Find the constant of variation for the relationship f(x)= 30x. A. 10 B. 30 C. x D. f(x) 2. If f(x) varies directly with x, and f(x) = 48 when x = 8, find the value of f(x) when x = 2 A. 12 B. 6 C. 8 D. 3

OpenStudy (anonymous):

Ok the constant of variation is a constant number that f(x) varies by Let say x=0 then f(x)=0 Lets say x=1 then f(x)=30 Lets say x=2 then f(x)=60 So f(x) is jumping by 30's so what would you assume is the constant of variation?

OpenStudy (anonymous):

so lost

OpenStudy (anonymous):

You so dont know the basics of algebra ... Ok lets start from the beggining

OpenStudy (anonymous):

We have an equation f(x)=30x We plug in values such as 1,2 ,3 ,4 5 and 6 .... into our x and then we find out what f(x) equals So let x=1 what will f(1) equal??? \( f(1)=30(1) = 30\) so f(1)=30 So when x - 1 we know that f(1)=30 Let x=2 what will f(2) equal ????? \( f(2)=30(2) = 60\) so f(2)=60 So we see that we plug in our x values to find out what f(x) will equal (Another term for f(x) is y) So our f(x) is dependent on what x is. We can only find out what f(x) equals once we know what our x equals.

OpenStudy (anonymous):

@Alez do you follow so far?

OpenStudy (anonymous):

no? lol

OpenStudy (anonymous):

srry have been depressed alot and not enough sleef so ya its hard for me to understand alot of things

OpenStudy (anonymous):

@johnweldon1993 Can you help out here. Not sure how else to explain this topic

OpenStudy (anonymous):

What is a constant?? Its a number that doesnt change In that equation x can take on any value. We can make x be 1 or 2 or 4 or any number we want and then f(x) equals a number depending on what x equals so both f(x) and x are not constant. They can be many values. Only the 30 stays the same. It is constant